How to sove this
(4/2-1/3)+x=13/15

Answer by checkley71(8403)

(4/2-1/3)+X=13/15 OR 2-1/3+X=13/15 OR 1 2/3+X=13/15 OR X=13/15-1 2/3 OR
X=13/15-5/3 OR X=13/15-25/15 OR X=-12/15